OpenStudy (song_of_the_sole):

A total of 20.0 mL of sodiutm hydroxide (NaOH) was neutalized by 30.0 mL of 0.250 M hydrogen bromide (HBr). What was the concentration of the sodium hydroxide? ****** I need help filling out the chart. please help me through it

OpenStudy (kyanthedoodle):

@imqwerty

imqwerty (imqwerty):

The balanced reaction of the two chemicals will be this- \(NaOH + HBr \) ----> \(NaBr+H_2O\) here \(\color{orangered}{1 ~mole~ of~ HBr~neutralizes~ 1 ~mole~ of ~NaOH}\) so we are given all this info- the concentration of \(HBr\) =0.25M volume of \(HBr\) = 30mL and volume of \(NaOH\)=20mL and it says that A total of 20.0 mL of sodiutm hydroxide (NaOH) was neutalized by 30.0 mL of 0.250 M hydrogen bromide (HBr) ^here the number of moles of \(NaOH\) will be equal to the number of moles of \(HBr\) because \(HBr\) is completely neutralizing the \(NaOH\) so we find the number of moles of \(HBr\) with the given info \(\large Molarity =\frac{moles_{of~HBr}}{volume_{in ~mL}} \times 1000\) \(0.25= \frac{moles}{30} \times 1000 \) \(moles= \frac{0.25 \times 30}{1000}\) okay so we have the number of moles of \(HBr\) the number of moles of \(NaOH\) will also be this much so \(Molarity= \frac{moles}{volume_{in~mL} } \times 1000\) \(Molarity= \frac{\frac{0.25 \times 30}{1000}}{20} \times 1000\) \(Molarity= \frac{0.25 \times 30}{1000 \times 20} \times 1000\) \(Molarity= \frac{0.25 \times 30}{20}\) \(Molarity=0.375M\)

OpenStudy (song_of_the_sole):

im a little confused on the molar ratio would it be 1 for HBr and 1 for NaOH?

imqwerty (imqwerty):

you mean why will be the moles equal? :)

OpenStudy (song_of_the_sole):

yes

imqwerty (imqwerty):

okay see the reaction- \(\color{orangered}{1}HBr+\color{orangered}{1}NaOH\) -----------> \(\color{orangered}{1}NaBr+\color{orangered}{1}H_2O\) ^this is a balanced stoichiometric reaction and in a balanced stoichiometric reaction the stoichiometric coefficients[the numbers in the red :) ] of each chemical denote the ratio of the moles of the chemicals reacting or being produced. so here 1 mole of \(NaOH\) reacts with 1 mole of \(HBr\) and neutralizes completely so since the number of moles reacting to complete the neutralization reaction are same for both we equaled them suppose if 5 moles of \(NaOH\) react then correspondingly 5 moles of \(HBr\) will be needed to complete the neutralization reaction :)

OpenStudy (song_of_the_sole):

oh okay that makes sense.

imqwerty (imqwerty):

:)

OpenStudy (song_of_the_sole):

how do i find the concentration of NaOH now?

imqwerty (imqwerty):

concentration is nothing but the molarity molarity is given by this expression- \(molarity= \frac{moles}{Volume} \times 1000\) here the volume must be taken in mL. alright so if we wish to find the molarity of \(NaOH\) - we know the volume but we donno the moles so we can't find it rn but we know that since the NaOH is getting completely neutralized its moles must be equal to the moles of HBr so next step is to find the moles of HBr

OpenStudy (song_of_the_sole):

and how do we do that?

imqwerty (imqwerty):

we know the molarity and the volume of HBr try to put these values into that molarity equation to get the number of moles :)

OpenStudy (song_of_the_sole):

im not sure what numbers i need to put in for the equation

OpenStudy (song_of_the_sole):

sorry i really suck at chemistry

imqwerty (imqwerty):

okay this is our equation- \(molarity = \frac{moles}{volume} \times 1000\) read the question its given that the molarity of HBr is \(0.25M\) and the volume of HBr is \(30mL\) we can plug this info into that equation to get moles?

OpenStudy (song_of_the_sole):